OpenStudy (anonymous):

I need to graph the equation and indetify the y-interception y=1/2 x

OpenStudy (anonymous):

To find the y intercept algebraically, which is in my opinion much easier for an equation like this, set y=0 and solve for x. If you must do it graphically for an assignment there are a variety of online tools that will graph simple equations for you. Try wolfram alpha for example: http://www.wolframalpha.com/input/?i=y%3D1%2F2+x

OpenStudy (anonymous):

Thank you...I am having trouble with the y intercept. the answer that I have for teh graphing is y=x/2 . I will try the wolframalpha website.

OpenStudy (anonymous):

if y=1/2x that means that the y intercept will be 0

OpenStudy (callisto):

@karinalol well... yes

OpenStudy (anonymous):

The ordered pair is (0,2)?

OpenStudy (callisto):

@stilenx for y-int, it is at x=0, so you should put x=0 into the equation instead of y=0

OpenStudy (callisto):



OpenStudy (callisto):

and if you have the equation in the form y=mx+c, like the one in the question, y=(1/2)x m= slope , and c= y-intercept so from the equation, you can directly find out that the slope is 1/2 and the y-int is 0 (since you have no numbers behind 'mx')

OpenStudy (anonymous):

I think I have that part of the question? I need to find the ordered pair?

OpenStudy (callisto):

not sure, but to graph the equation, you need to find 2 points

OpenStudy (anonymous):

I graphed the equation correctly and now I have to provide an ordered pa

OpenStudy (callisto):

what order pair?

OpenStudy (anonymous):